Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by Jacwis » Sun Apr 21, 2019 6:09 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Hi

I need help with the following question.

Four people each will roll a fair die once. What is the probability that at least two will roll the same number?

A) 50%
B) 65%
C) 80%
D) 72%
E) 40%

The problem can easily be solved by making all the people roll the same number then subtracting the result from 1. However, I wanted to understand the problem by adding up all the cases i.e two roll the same number or three roll the same number etc. but i am not getting the correct answer. Can any expert help me find the mistake?


Case 1 Two people throw the same number and the other two each roll a different number. e.g 4456

6/6 * 1/6 *5/6 * 4/6 *4!/3!2! I am not sure how to calculate all the permutations of this. To arrange four numbers i multiplied by 4! then divided by 3! to eliminate double counting and again by 2! to remove the same digit cases i.e 44 55.

Case 2 two roll the same number and the other two roll the same but different from the one that was rolled by the first two people. e.g 3344

6/6* 1/6* 5/6* 1/6 *4/2! 2! 2! Here I arrange the four numbers then divided by 2! three times to remove double digits already arranged in the product and to eliminate two same digit groups. again i am not sure if this is correct way to count all permutations of this case.


Case 3 three roll the same number and the third rolls a different number eg 4445

6/6 * 1/6*1/6*5/6 * 4!/ 2!*3!


Case 4 All roll the same number

this is easy 6/6^4

Adding up all the cases does not yield the right answer. Please help me find the mistake. Thanks

Jac
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Mon Apr 22, 2019 3:45 am
Jacwis wrote:Hi

I need help with the following question.

Four people each will roll a fair die once. What is the probability that at least two will roll the same number?

A) 50%
B) 65%
C) 80%
D) 72%
E) 40%

I wanted to understand the problem by adding up all the cases
All possible rolls:
Since there are 6 number options for each roll, we get:
6*6*6*6 = 1296

Case 1: Exactly 2 rolls are the same (such as 1123)
From 4 rolls, the number of ways to choose a pair to yield the same number = 4C2 = (4*3)/(2*1) = 6.
Number options for this pair = 6. (Any of the 6 numbers on the die)
Number options for the third roll = 5. (Any of the 5 remaining numbers on the die)
Number options for the last roll = 4. (Any of the 4 remaining numbers on the die).
To combine these options, we multiply:
6*6*5*4 = 720

Case 2: Two pairs (such as 1122)
Number of ways to divide 4 rolls into 2 pairs = 3.
(First way: 1st and 2nd rolls are the same; 3rd and 4th rolls are the same but different from the 1st and 2nd.)
(Second way: 1st and 3rd rolls are the same; 2nd and 4th rolls are the same but different from the 1st and 3rd.)
(Third way: 1st and 4th rolls are the same; 2nd and 3rd rolls are the same but different from the 1st and 4th.)

Number options for first pair = 6. (Any of the 6 numbers on the die)
Number options for the second pair = 5. (Any of the 5 remaining numbers on the die)
To combine these options, we multiply:
3*6*5 = 90

Case 3: Exactly 3 numbers are the same
From 4 rolls, the number of ways to choose three to yield the same number = 4C3 = (4*3*2)/(3*2*1) = 4.
Number options for these 3 rolls = 6. (Any of the 6 numbers on the die)
Number options for the fourth roll = 5. (Any of the 5 remaining numbers on the die)
To combine these options, we multiply:
4*6*5 = 120

Case 4: All 4 numbers are the same
Number options = 6. (Any of the 6 numbers on the die)

Case 1 + Case 2 + Case 3 + Case 4 = 720 + 90 + 120 + 6 = 936
Thus:
P(at least 2 people roll the same number) = 936/1296 ≈ 0.72 = 72%

The correct answer is D.

MUCH easier to solve as follows:

P(at least 2 numbers are the same) = 1 - P(all 4 numbers are different)

The first person can roll any of the 6 numbers.
P(2nd number is different from the 1st) = 5/6
P(3rd number is different from the first 2 numbers) = 4/6
P(4th number is different from the first 3 numbers) = 3/6
To combine these probabilities, we multiply:
5/6 * 4/6 * 3/6 = 5/6 * 2/3 * 1/2 = 5/18

Thus:
P(at least 2 numbers are the same) = 1 - 5/18 = 13/18 ≈ 0.72 = 72%
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Jacwis » Thu Apr 25, 2019 12:24 am
Thanks Mitch for the solution

I have some confusion, in the second case in your post, 1122 . When we select two positions for a pair from the four positions, it can be done in 4C2=6 ways. However, when we can place any pair in 6 ways then the second pair can be placed in the remaining two positions in only 1 way. Isn't it so?Why are there 3 ways to roll the two pairs? For example 1122 can be arranged 4!/2!2! which is 6.
All the cases of 1122 are
1122
1212
1221
2211
2121
2112

What is wrong with this reasoning. Please, elaborate. Thanks
Join the discussion

probability

by GMATGuruNY » Thu Apr 25, 2019 2:54 am
Jacwis wrote:two roll the same number and the other two roll the same but different from the one that was rolled by the first two people. e.g 3344

6/6* 1/6* 5/6* 1/6 *4/2! 2! 2! Here I arrange the four numbers then divided by 2! three times to remove double digits already arranged in the product and to eliminate two same digit groups. again i am not sure if this is correct way to count all permutations of this case.

When we select two positions for a pair from the four positions, it can be done in 4C2=6 ways. However, when we can place any pair in 6 ways then the second pair can be placed in the remaining two positions in only 1 way. Isn't it so?Why are there 3 ways to roll the two pairs? For example 1122 can be arranged 4!/2!2! which is 6.
All the cases of 1122 are
1122
1212
1221
2211
2121
2112

What is wrong with this reasoning
The approach above counts the following:

Cases that include 11:
1122, 1133, 1144, 1155, 1166
Cases that include 22:
2211, 2233, 2244, 2255, 2266
Cases that include 33:
3311, 3322, 3344, 3355, 3366
Cases that include 44:
4411, 4422, 4433, 4455, 4466
Cases that include 55:
5511, 5522, 5533, 5544, 5566
Cases that include 66:
6611, 6622, 6633, 6644, 6655

Notice the following:
The cases in red have all been double-counted.
1122 will yield the same outcomes as 2211.
1133 will yield the same outcomes as 3311.
1144 will yield the same outcomes as 4411.
1155 will yield the same outcomes as 5511.
1166 will yield the same outcomes as 6611.
By extension, EVERY case above has been double-counted.
To account for the double-counting, we must divide by 2:
(6/6* 1/6* 5/6* 1/6)(4!/2!2!)/2
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion